OpenStudy (anonymous):

Are the roots of the equation x^2-10x+25=0 imaginary?

OpenStudy (nathan917):

1. x^2 - 10x + 25 = 0 b^2 - 4ac = (-1)^2 - 4(1)(25) = 100 - 100 = 0 c

OpenStudy (anonymous):

The discriminant tells us what type of roots a quadratic has. It has imaginary roots if \[b^2 - 4ac < 0.\]

OpenStudy (anonymous):

is it real, rational, irrational, equal,unequal ?

OpenStudy (anonymous):

The discriminant is zero so we have repeated real roots.